# RRCANet: Recurrent Reusable-Convolution Attention Network for Infrared Small Target Detection

### v2

@@ Abstract
Infrared small target detection is a challenging task due to its unique characteristics (e.g., small, dim, shapeless and changeable). Recently published CNN-based methods have achieved promising performance with heavy feature extraction and fusion modules. To achieve efficient and effective detection, we propose a recurrent reusable-convolution attention network (RRCA-Net) for infrared small target detection. Specifically, RRCA-Net incorporates reusable-convolution block (RuCB) in a recurrent manner without introducing extra parameters. With the help of the repetitive iteration in RuCB, the high-level information of small targets in the deep layers can be well maintained and further refined. Then, a dual interactive attention aggregation module (DIAAM) is proposed to promote the mutual enhancement and fusion of refined information. In this way, RRCA-Net can both achieve high-level feature refinement and enhance the correlation of contextual information between adjacent layers. Moreover, to achieve steady convergence, we design a target characteristic inspired loss function (DpT-k loss) by integrating physical and mathematical constraints. Experimental results on three benchmark datasets (e.g. NUAA-SIRST, IRSTD-1k, DenseSIRST) demonstrate that our RRCA-Net can achieve comparable performance to the state-of-the-art methods while maintaining a small number of parameters, and act as a plug and play module to introduce consistent performance improvement for several popular IRSTD methods. Our code will be available at https://github.com/yongxianLiu/RRCANet soon.
Index Terms:
Infrared small target detection, reusable convolution block, dual attention, characteristic inspired loss function.
@@ I Introduction
Figure 1: A comparison between our RRCANet and recently published methods in term of IoU, FLOPs and number of parameters. Larger circle represents more parameters.
Deep-learning-based methods can interpret the high-level semantic correlation between small targets and clutter background with a data-driven manner and thus achieve high-precision detection. Dai et al. [13] proposed an asymmetric contextual module (ACM) to combine top-down global context feedback with a bottom-up modulated path for highlighting small targets. Then, Li et al. [14] designed the dense nested interaction module (DNIM) and cascade channel and spatial attention module (CSAM) to preserve small target information and enhance multi-scale features. Wu et al. [15] embedded a tiny U-Net into a larger U-Net backbone to form nested structure, which effectively addresses the problem that feature resolution decreases with the increase of network depth in neural networks. Next, ISNet [16] extracted edge and shape features to reconstruct targets, effectively improving the detection accuracy. Recently, Yuan et al. [17] proposed a spatial-channel cross transformer block (SCTBs) to enlarge the semantic differences between targets and clutter at all levels. Although these methods have achieved promising performance, they mainly rely on highly stacked network structure, dense connections and heavy multi-scale feature fusion. A simple and lightweight detection network is expected.
Figure 2: The representation of small targets in CNN with recurrent iteration of (a) RRCA-Net without iteration (b) RRCA-Net with n=2.
@@
  • A resuable-convolution block (RuCB) and a dual interactive attention aggregation module (DIAAM) are proposed to achieve high-level feature refinement and enhance the contextual information aggregation capability.
  • We develop a small target characteristic inspired DpT-k loss to facilitate network convergence by introducing physical and mathematical constraints. Experimental results on three benchmark datasets (e.g. NUAA-SIRST, IRSTD-1k and DenseSIRST) show that our methods achieve comparable detection performance with fewer parameters. (as shown in Fig. 1).
This paper is organized as follows. Section II briefly reviews infrared small target detection methods, lightweight detection networks, and specially designed loss functions. Section III introduces the architecture of our RRCA-Net and our DpT-k loss. Experimental results on three single-frame dataset are discussed in Section IV. The last Section V shows the conclusion.
@@ II Related Work
In this section, we mainly review the main work on infrared small target detection (IRSTD), lightweight detection networks and specially designed loss functions for IRSTD.
@@ II-A Infrared Small Target Detection
Deep-learning-based methods can learn robust representation via a data-driven manner. Reachers have introduced a variety of neural networks to IRSTD. Many detection methods [13, 14, 15] were based on CNN and designed many target characteristic-inspired modules to improve detection performance, including spatial attention, dense connection, module nesting and so on. Highly stacked modules and heavy feature fusion strategies significantly improved the detection performance but introduced huge computation burden. However, CNN-based methods lack the long-range feature aggregation ability. To handle this problem, MTU-Net [19] applied a multi-level Vision Transformer (ViT)-CNN hybrid encoder to extract and integrate local details and global information. SCTransNet [17] leveraged SCTBs on skip connections to model effectually long-range information. It encoded the semantic difference between targets and background to enhance feature representation. Considering the high computational complexity, Chen et al. [20] proposed the nested Mamba network (MiM-ISTD) to achieve fast and low-memory detection. Besides, several new paradigms of neural networks have also been introduced into IRSTD. Jia et al. [21] proposed the IRGraphSeg based on graph deep learning, which can extract both image texture and structural information. Du et al. [22] treated IRSTD as a mask-generation task and designed a conditional diffusion framework.
@@
Figure 3: An illustration of the proposed reusable-convolution attention network (RRCA-Net). Feature extraction: Images are first input into the reusable-convolution block (RuCB) of encoder to extract multi-layer features. Feature aggregation and enhancement: Subsequently, in the decoder, the extracted features are upsampled and fused with a dual interactive attention aggregation module (DIAAM). Repeated feature refinement: The multi-layer features are concatenated to achieve robust output refinement after several encoding and decoding iterations. Finally, eight-connected neighborhood clustering algorithm clusters the segmentation map to locate the centroid of each target region.
@@ II-B Lightweight Networks for IRSTD
The first kind of lightweight method is model compression. These methods can be divided into the following categories: 1) Knowledge distillation [23, 24], which improves the performance of compact student model based on the knowledge of the cumbersome teacher networks. 2) Quantization [25], which uses lower bitwidth to represent network weight and activation. 3) Pruning [26, 27], which focuses on searching redundancy structure and then eliminating unimportant weights and channels. 4) Network architecture search [28, 29], which automatically explores the optimal network structure or component structure in the search space, combining with constraints and search strategies. 5) Low-rank factorization [30], which uses the low-rank constraints of network to decrease the computation.
The second kind of method focuses on designing lightweight network structure [31, 32, 33, 34, 35], including 1) reducing backbone depth, width, kernel size and improving accuracy with simple attention. 2) replacing traditional convolution with lightweight convolution operators, such as separable convolution, group convolution and sparse convolution. 3) accelerating model inference through re-parameterization. These methods heavily rely on manual design and are therefore not as general as model compression methods. Although the computational cost is further reduced, existing lightweight methods require lots of empirical knowledge and manual selection. More simple yet effective lightweight detection methods are expected to be explored.
@@ II-C Loss Functions for IRSTD
Appropriate loss function can accelerate the network convergence and increase the detection performance. Existing loss functions of small target detection field can be divided into three categories: 1) Distribution-based loss aims to minimize the distribution difference between predictions and ground truth, including cross-entropy loss [36], Focal loss [37]. 2) Region-based loss aims to maximize the overlap area between predictions and masks, including Dice loss [38], SoftIoU loss [39]. 3) Composite loss combines multiple types of losses, each of which complements the strengths of the others, including weighting, power and exponential. Considering that IRSTD has highly unbalanced positive and negative samples, researches have designed many composite loss functions to achieve better and quicker network convergence. Zhang et al. [16] used Dice loss and Edge loss to achieve accurate prediction of the target edge and shape. Edge loss (i.e., BCE loss + Dice loss) was designed to strengthen the utilization of target edge information. In space-based ship detection, Wu et al. [19] combined the advantages of Focal loss and SoftIoU loss to strengthen target localization and shape description. Focusing on lack of sensitivity to target scale and position, Liu et al. [40] developed a novel scale and location sensitive (SLS) loss, enabling existing detectors to improve detection performance.
@@ III Methodology
In this section, we introduce the RRCA-Net, RuCB block (Section III-B), DIAAM (Section III-C) and DpT-k loss (Section III-F) in detail.
@@ III-B2 The Reusable-Convolution Block in Encoder
where $\cal{R}(\cdot)$ denotes the cascaded reusable-convolution block.
III-C The Simplified Asymptotic Decoder
III-C1 The Dual Interactive Attention Aggregation Module
@@ III-D The Feature Stacking Module
[…]
@@ III-E The Eight-connected Domain Clustering Algorithm for Evaluation
After obtaining the output feature map, we introduce an eight-connected domain clustering algorithm [14] to calculate the evaluation metrics(i.e. $P_{d}$ and $F_{a}$). In this way, we can cluster the pixels within the same connected domain and then compute the centroid coordinates of each target. In feature maps $\mathbf{G}$, when a pixel ${(x_{0},y_{0})}$ is in the eight-neighborhood of another pixel ${(x_{1},y_{1})}$ and these two pixels have the same value (0 or 1), they are judged to belong to the same connected domain, i.e.,
@@ III-F DpT-k Loss for Infrared Small Target Segmentation
Top-k loss retains top $k\%$ of the worst predicted pixel for calculating values, which aims to force network to focus on hard samples and adapts to imbalanced distributions. It is defined by:
@@ IV Experiment
In this section, we first introduce our evaluation metrics and implementation details of comparative experiments. Then, we compare our RRCA-Net to several SOTA IRSTD methods and lightweight networks. Finally, we present ablation studies to investigate the intrinsic reasons why our network works.
@@ IV-A Evaluation Metrics
Detection Performance Metrics: Infrared small targets generally have less pixels and are lack of textures and shapes. Because of tiny target area, the unsuitable evaluation metrics can easily cause misjudgment. Following the pioneering works based on segmentation [14, 19], we adopt pixel-level criteria to evaluate performance. Probability of detection (${P}_{d}$) and false alarm rate (${F}_{a}$) are used to evaluate the localization competency. Intersection over union (IoU) is used to evaluate the shape description competency. In addition, the receiver operating characteristic (ROC) curves can comprehensively reflect the overall performance of the network on ${P}_{d}$ and ${F}_{a}$.
Computation Cost Metrics: The computational complexity analysis is important to measure model size and computational cost when deploying model on resource-limited devices. ##\##Params is used to count the number of learnable parameters in neural networks and can also reflect model size.
TABLE I: Results comparison of detection performance ($IoU$(×$\times 10^{-2}$), $P_{d}$(×$\times 10^{-2}$), and $F_{a}$(×$\times 10^{-6}$)) and model size (#$\#Params$(M)) achieved by different state-of-the-art methods on three datasets. The best results are in red and the second best results are in blue.
@@
Top-Hat [5] 10.06 75.11 1432 7.14 79.84 1012 1.07 32.77 563.9 -
Max-Median [6] 7.00 65.21 59.73 4.17 69.20 55.33 0.12 33.24 5.178 -
WSLCM [7] 3.45 72.44 6619 1.16 77.95 5446 2.58 48.90 2100 -
TLLCM [8] 3.31 77.39 6738 1.03 79.09 5899 1.12 22.15 2300 -
IPI [9] 27.92 81.37 16.18 25.67 85.55 11.47 3.35 42.13 10.45 -
NRAM [10] 15.25 70.68 16.93 12.16 74.52 13.85 4.91 65.32 4.363 -
RIPT [11] 14.11 77.55 28.31 11.05 79.08 22.61 5.98 72.93 9.179 -
PSTNN [12] 24.57 71.99 35.26 22.40 77.95 29.11 4.77 35.12 138.7 -
MSLSTIPT [4] 11.43 79.03 1524 10.30 82.13 1131 2.13 17.37 333.2 -
ACM [13] 60.33 92.27 65.31 68.34 91.63 23.19 53.64 61.75 31.56 0.398
ResUNet [45] 66.06 90.48 10.63 74.06 96.96 14.97 80.96 98.06 4.377 0.9141
AMFU-net [46] 66.70 92.86 16.02 74.26 97.34 14.40 80.33 97.89 4.132 0.4730
DNANet [14] 67.70 91.58 18.69 76.63 95.44 26.55 81.45 97.28 3.252 4.697
UIU-Net [15] 65.69 91.25 13.42 76.24 95.06 10.57 77.51 97.11 6.699 50.54
LW-IRSTNet [33] 61.34 89.46 23.84 67.83 94.30 19.39 78.95 97.11 3.656 0.163
RepISD-Net [34] 65.45 91.59 7.62 74.87 98.48 6.85 78.60 97.25 22.24 0.279
RRCA-Net-ResNet10 (ours) 66.57 90.48 13.36 74.17 96.20 20.25 82.35 98.03 1.319 0.103
RRCA-Net-ResNet18 (ours) 68.73 92.52 6.68 75.41 98.10 7.56 82.15 98.23 1.810 0.202
RRCA-Net-ResNet34 (ours) 67.31 92.52 7.29 72.45 96.96 12.34 82.00 98.49 2.308 0.363
@@ IV-A2 Probability of Detection
${P}_{d}=\frac{Target_{correct}}{Target_{All}}.$ (12)

### v3

@@ Abstract
Infrared small target detection is a challenging task due to its unique characteristics (e.g., small, dim, shapeless and changeable). Recently published CNN-based methods have achieved promising performance with heavy feature extraction and fusion modules. To achieve efficient and effective detection, we propose a recurrent reusable-convolution attention network (RRCA-Net) for infrared small target detection. Specifically, RRCA-Net incorporates reusable-convolution block (RuCB) in a recurrent manner without introducing extra parameters. With the help of the repetitive iteration in RuCB, the high-level information of small targets in the deep layers can be well maintained and further refined. Then, a dual interactive attention aggregation module (DIAAM) is proposed to promote the mutual enhancement and fusion of refined information. In this way, RRCA-Net can both achieve high-level feature refinement and enhance the correlation of contextual information between adjacent layers. Moreover, to achieve steady convergence, we design a target characteristic inspired loss function (DpT-k loss) by integrating physical and mathematical constraints. Experimental results on three benchmark datasets (e.g. NUAA-SIRST, IRSTD-1k, DenseSIRST) demonstrate that our RRCA-Net can achieve comparable performance to the state-of-the-art methods while maintaining a small number of parameters, and act as a plug and play module to introduce consistent performance improvement for several popular IRSTD methods.
@@ I Introduction
Figure 1: Comparison between our RRCANet and recently published methods in term of IoU, FLOPs and number of parameters. Larger circle represents more parameters.
Deep-learning-based methods can interpret the high-level semantic correlation between small targets and clutter background with a data-driven manner and thus achieve high-precision detection. Dai et al. [13] proposed an asymmetric contextual module (ACM) to combine top-down global context feedback with a bottom-up modulated path for highlighting small targets. Then, Li et al. [14] designed the dense nested interaction module (DNIM) and cascade channel and spatial attention module (CSAM) to preserve small target information and enhance multi-scale features. Wu et al. [15] embedded a tiny U-Net into a larger U-Net backbone to form nested structure, which effectively addresses the problem that feature resolution decreases with the increase of network depth in neural networks. Next, ISNet [16] extracted edge and shape features to reconstruct targets, effectively improving the detection accuracy. Recently, Yuan et al. [17] proposed a spatial-channel cross transformer block (SCTBs) to enlarge the semantic differences between targets and clutter at all levels. Although these methods have achieved promising performance, they mainly rely on highly stacked network structure, dense connections and heavy multi-scale feature fusion. Different from previous methods with massive modules and heavy computation burden, we aim to design a more lightweight manner to achieve feature refinement. A recurrent reusable convolution network is proposed to increase the detection performance but maintain comparable parameters with previous method.
Figure 2: Representation of small targets in CNN with recurrent iteration of (a) RRCA-Net without iteration (b) RRCA-Net with n=2.
@@
  • 2)
    An RuCB and a DIAAM are proposed to achieve high-level feature refinement and enhance the contextual information aggregation capability.
  • 3)
    We develop a small target characteristic inspired DpT-k loss to facilitate network convergence by introducing physical and mathematical constraints. Experimental results on three benchmark datasets (e.g. NUAA-SIRST, IRSTD-1k and DenseSIRST) show that our methods achieve comparable detection performance with fewer parameters. (see Fig. 1).
The rest of this article is organized as follows. Section II briefly reviews infrared small target detection (IRSTD) methods, lightweight detection networks, and specially designed loss functions. Section III introduces the architecture of our RRCA-Net and our DpT-k loss. Experimental results on three single-frame dataset are discussed in Section IV. Finally, Section V concludes this article.
@@ II Related Work
In this section, we mainly review the main work on IRSTD.
@@ II-A Infrared Small Target Detection
Deep-learning-based methods can learn robust representation via a data-driven manner. Reachers have introduced a variety of neural networks to IRSTD. Many detection methods [13, 14, 15] were based on CNN and designed many target characteristic-inspired modules to improve detection performance, including spatial attention, dense connection, module nesting and so on [19, 20]. Highly stacked modules and heavy feature fusion strategies significantly improved the detection performance but introduced huge computation burden. However, CNN-based methods lack the long-range feature aggregation ability. To handle this problem, MTU-Net [21] applied a multi-level Vision Transformer (ViT)-CNN hybrid encoder to extract and integrate local details and global information. Based on neural ordinary differential equations, RKformer [22] implemented the Runge-Kutta transformer blocks, which can effectively suppress noise. SCTransNet [17] leveraged SCTBs on skip connections to model effectually long-range information. It encoded the semantic difference between targets and background to enhance feature representation. Considering the high computational complexity, Chen et al. [23] proposed the nested Mamba network (MiM-ISTD) to achieve fast and low-memory detection. Besides, several new paradigms of neural networks have also been introduced into IRSTD. Zhang et al. [24] redesigned the segment anything model (SAM) for IRSTD task, demonstrating excellent performance. Then, Xu et al. [25] proposed SAMamba to efficiently deal with domain difference, which integrates hierarchical learning in SAM and global context modeling in Mamba. Jia et al. [26] proposed the IRGraphSeg based on graph deep learning, which can extract both image texture and structural information. Du et al. [27] treated IRSTD as a mask-generation task and designed a conditional diffusion framework. Shi et al.[28] achieved omni-domain infrared small target detection by incorporating auxiliary metadata.
@@ II-B Lightweight Networks for IRSTD
The first kind of lightweight method is model compression. These methods can be divided into the following categories.
  • 1)
    Knowledge distillation [29, 30], which improves the performance of compact student model based on the knowledge of the cumbersome teacher networks.
  • 2)
    Quantization [31], which uses lower bitwidth to represent network weight and activation.
  • 3)
    Pruning [32, 33], which focuses on searching redundancy structure and then eliminating unimportant weights and channels.
  • 4)
    Network architecture search [34, 35], which automatically explores the optimal network structure or component structure in the search space, combining with constraints and search strategies.
  • 5)
    Low-rank factorization [36], which uses the low-rank constraints of network to decrease the computation.
The second kind of method focuses on designing lightweight network structure [37, 38, 39, 40, 41], including the following.
  • 1)
    Reducing backbone depth, width, kernel size and improving accuracy with simple attention.
  • 2)
    Replacing traditional convolution with lightweight convolution operators, such as separable convolution, group convolution and sparse convolution.
  • 3)
    Accelerating model inference through re-parameterization. These methods heavily rely on manual design and are therefore not as general as model compression methods.
Although the computational cost is further reduced, existing lightweight methods require lots of empirical knowledge and manual selection. More simple yet effective lightweight detection methods are expected to be explored.
@@ II-C Loss Functions for IRSTD
Appropriate loss function can accelerate the network convergence and increase the detection performance. Existing loss functions of small target detection field can be divided into three categories as follows.
  • 1)
    Distribution-based loss aims to minimize the distribution difference between predictions and ground truth, including cross-entropy loss [42], Focal loss [43].
  • 2)
    Region-based loss aims to maximize the overlap area between predictions and masks, including Dice loss [44], SoftIoU loss [45].
  • 3)
    Composite loss combines multiple types of losses, each of which complements the strengths of the others, including weighting, power and exponential.
Considering that IRSTD has highly unbalanced positive and negative samples, researches have designed many composite loss functions to achieve better and quicker network convergence. Zhang et al. [16] used Dice loss and Edge loss to achieve accurate prediction of the target edge and shape. Edge loss (i.e., BCE loss + Dice loss) was designed to strengthen the utilization of target edge information. In space-based ship detection, Wu et al. [21] combined the advantages of Focal loss and SoftIoU loss to strengthen target localization and shape description. Chen et al.[46] combined Dice loss and Focal loss to jointly optimize detection and segmentation tasks. Focusing on lack of sensitivity to target scale and position, Liu et al. [47] developed a novel scale and location sensitive (SLS) loss, enabling existing detectors to improve detection performance.
Figure 3: An illustration of the proposed RRCA-Net. Feature extraction: Images are first input into the RuCB of encoder to extract multi-layer features. Feature aggregation and enhancement: Subsequently, in the decoder, the extracted features are upsampled and fused with a DIAAM. Repeated feature refinement: The multi-layer features are concatenated to achieve robust output refinement after several encoding and decoding iterations. Finally, eight-connected neighborhood clustering algorithm clusters the segmentation map to locate the centroid of each target region.
@@ III Methodology
In this section, we introduce the RRCA-Net and its components in detail.
@@ III-B2 The Reusable-Convolution Block in Encoder
[…]
III-C The Dual Interactive Attention Aggregation Module
@@ III-D The Feature Stacking Module
where $\left[\cdot,\cdot\right]$ represents the concatenate operation.
@@ III-E The Eight-connected Domain Clustering Algorithm for Evaluation
After obtaining the output, we introduce an eight-connected domain clustering algorithm to calculate $P_{d}$ and $F_{a}$. In this way, we can cluster the pixels within the same connected domain and then compute the centroid coordinates of each target. In feature maps $\mathbf{G}$, when a pixel $(x_{0},y_{0})$} is in the eight-neighborhood of another pixel $(x_{1},y_{1})$} and these two pixels have the same value (0 or 1), they are judged to belong to the same connected domain, i.e.,
@@ III-F DpT-k Loss for Infrared Small Target Segmentation
Top-k loss retains top $k\%$ of the worst predicted pixel for calculating values, which aims to force network to focus on hard samples and adapts to imbalanced distributions. In other words, the ranking criterion is the value of cross entropy between predicted pixels and groundtruth pixels. The value of Top-k is calculated by the largest k% values of all cross entropy between pixels. It is defined by:
@@ IV Experiment
In this section, we first introduce our evaluation metrics and implementation details. Then, comparison experiments and ablation studies are presented.
@@ IV-A Evaluation Metrics
Detection performance metrics include intersection over union (IoU), probability of detection (${P}_{d}$), false alarm rate (${F}_{a}$) and receiver operating characteristic (ROC) curves.
Computation cost metrics include #\#Params and FLOPs.
TABLE I: Results comparison of detection performance (IoUIoU($\times 10^{-2}$), $P_{d}$($\times 10^{-2}$), and $F_{a}$($\times 10^{-6}$)) and model size (#Params\#Params(M)) achieved by different state-of-the-art methods on three datasets. The best results are in red and the second best results are in blue.
@@
Top-Hat [5] 10.06 75.11 1432 7.14 79.84 1012 1.07 32.77 563.9 - -
Max-Median [6] 7.00 65.21 59.73 4.17 69.20 55.33 0.12 33.24 5.178 - -
WSLCM [7] 3.45 72.44 6619 1.16 77.95 5446 2.58 48.90 2100 - -
TLLCM [8] 3.31 77.39 6738 1.03 79.09 5899 1.12 22.15 2300 - -
IPI [9] 27.92 81.37 16.18 25.67 85.55 11.47 3.35 42.13 10.45 - -
NRAM [10] 15.25 70.68 16.93 12.16 74.52 13.85 4.91 65.32 4.363 - -
RIPT [11] 14.11 77.55 28.31 11.05 79.08 22.61 5.98 72.93 9.179 - -
PSTNN [12] 24.57 71.99 35.26 22.40 77.95 29.11 4.77 35.12 138.7 - -
MSLSTIPT [4] 11.43 79.03 1524 10.30 82.13 1131 2.13 17.37 333.2 - -
ACM [13] 60.33 92.27 65.31 68.34 91.63 23.19 53.64 61.75 31.56 0.398 0.335
ResUNet [52] 66.06 90.48 10.63 74.06 96.96 14.97 80.96 98.06 4.377 0.914 1.27
AMFU-net [53] 66.70 92.86 16.02 74.26 97.34 14.40 80.33 97.89 4.132 0.473 5.09
DNANet [14] 67.70 91.58 18.69 76.63 95.44 26.55 81.45 97.28 3.252 4.697 14.05
UIU-Net [15] 65.69 91.25 13.42 76.24 95.06 10.57 77.51 97.11 6.699 50.54 46.65
LW-IRSTNet [39] 61.34 89.46 23.84 67.83 94.30 19.39 78.95 97.11 3.656 0.163 0.177
RepISD-Net [40] 65.45 91.59 7.62 74.87 98.48 6.85 78.60 97.25 22.24 0.279 11.53
RRCA-Net-ResNet10 (ours) 66.57 90.48 13.36 74.17 96.20 20.25 82.35 98.03 1.319 0.103 0.721
RRCA-Net-ResNet18 (ours) 68.73 92.52 6.68 75.41 98.10 7.56 82.15 98.23 1.810 0.202 1.04
RRCA-Net-ResNet34 (ours) 67.31 92.52 7.29 72.45 96.96 12.34 82.00 98.49 2.308 0.363 1.68
@@ IV-A2 Probability of Detection
${P}_{d}=\frac{Target_{correct}}{Target_{All}},$ (12)